OpenStudy (anonymous):

what is the square of 8???

OpenStudy (zzr0ck3r):

8 squared, or sqrt root of 8?

OpenStudy (zzr0ck3r):

sqrt(8) = sqrt(4*2) = sqrt(4)*sqrt(2) = 2*sqrt(2) 8^2 = 8*8 = 64

OpenStudy (anonymous):

thank u